Lenow’s Drug Stores and Hall’s Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here.

Lenow Hall
Debt @ 10%
$ 290,000
Debt @ 10%
$ 580,000
Common stock, $10 par 580,000 Common stock, $10 par 290,000
Total
$ 870,000
Total
$ 870,000
Common shares
58,000
Common shares
29,000

a. Complete the following table given earnings before interest and taxes of $33,000, $87,000, and $88,000. Assume the tax rate is 30 percent. (Negative amounts should be indicated by parentheses or a minus sign. Round your answers to 2 decimal places.)

What is the relation ship between
EBIT Total Assets EBIT/TA % Lenow EPS Hall EPS The EPS of the two firms
$33,000 $870,000 3.79 %
$87,000 $870,000 10.00 %
$88,000 $870,000 10.11 %

b-1. What is the EBIT/TA rate when the firm's have equal EPS?
EBIT/TA rate ____%

b-2. What is the cost of debt?
Cost of debt _____%

c. If the cost of debt went up to 12 percent and all other factors remained equal, what would be the break-even level for EBIT?
Break even level _____
Answer #1
a) EBIT Total Assets EBIT/TA % Lenow EPS Hall EPS What is the relationship between the EPS of the two firms
$         33,000 $      8,70,000 3.79 % $           0.05 $           -0.60 Lenow EPS>Hall EPS
$         87,000 $      8,70,000 10.00 % $           0.70 $             0.70 Leno EPS=Hall EPS
$         88,000 $      8,70,000 10.11 % $           0.71 $             0.72 Hall EPS>Lenow EPS
WORKSHEET:
LENOW:
EBIT Interest EBT Tax at 30% NI Number of shares EPS
$         33,000 $         29,000 $         4,000 $         1,200 $          2,800 58000 $         0.05
$         87,000 $         29,000 $       58,000 $      17,400 $        40,600 58000 $         0.70
$         88,000 $         29,000 $       59,000 $      17,700 $        41,300 58000 $         0.71
HALL:
$         33,000 $         58,000 $     -25,000 $       -7,500 $      -17,500 29000 $        -0.60
$         87,000 $         58,000 $       29,000 $         8,700 $        20,300 29000 $         0.70
$         88,000 $         58,000 $       30,000 $         9,000 $        21,000 29000 $         0.72
b-1) EBIT/TA rate = 10%.
b-2) Cost of debt = 10%
b-3) Break even level of EBIT is that level of EBIT at which EPS is 0 for the two firms.
EPS for Lenow = (E-34800)*0.7/58000 and
EPS for Hall = (E-69600)*0.7/29000
where E = the break even EBIT level.
Solving for E
(E-34800)*0.7/58000 = (E-69600)*0.7/29000
(E-34800) = (E-69600)*2
E - 34800 = 2*E - 139200
E = 139200-34800 = 104400
Break even level of EBIT = $104400.
CHECK:
EBIT Interest EBT Tax at 30% NI Number of shares EPS
$     1,04,400 $         34,800 $       69,600 $      20,880 $        48,720 58000 $         0.84
HALL:
$     1,04,400 $         69,600 $       34,800 $      10,440 $        24,360 29000 $         0.84